# Online Convex Optimization with Heavy Tails: Old Algorithms, New Regrets, and Applications

### v1

@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
$\mathbb{\mathbb{E}}\left[\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right],\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\right]\geq\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{t})-F(\boldsymbol{y}_{t})\right],$
where the last step is due to the convexity of $F$. As such, we can take expectations on both sides of (29) to have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{t})-F(\boldsymbol{y}_{t})\right]$ $\leq\frac{\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq\frac{\mathbb{\mathbb{E}}\left[\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}},$ (30)
where the second step is due to $\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\leq\left(1-\frac{w_{t-1}}{w_{t}}\right)\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t}\right\|^{2}+\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}=\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}$ by (26) and the convexity of $\left\|\boldsymbol{x}_{t}-\cdot\right\|^{2}$. Mutiply both sides of (30) by $w_{t}\eta_{t}$ and sum up from $t=1$ to $T$ to obtain
$\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}w_{t}\eta_{t}\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{y}_{t})\right)\right]$
$\leq$ $\frac{w_{0}\left\|\boldsymbol{x}_{1}-\boldsymbol{y}_{0}\right\|^{2}-\mathbb{\mathbb{E}}\left[w_{T}\left\|\boldsymbol{x}_{T+1}-\boldsymbol{y}_{T}\right\|^{2}\right]}{2}+\sum_{t=1}^{T}w_{t}\eta_{t}^{2}G^{2}+\mathsf{C}(\mathsf{p})w_{t}\eta_{t}^{\mathsf{p}}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq$ $\frac{w_{0}D^{2}}{2}+\sum_{t=1}^{T}w_{t}\eta_{t}^{2}G^{2}+\mathsf{C}(\mathsf{p})w_{t}\eta_{t}^{\mathsf{p}}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}.$ (31)
@@
$=\sum_{s=1}^{t}\frac{w_{s}-w_{s-1}}{w_{t}}\left(F(\boldsymbol{x}_{s})-F(\boldsymbol{x})\right),$
@@
$\sum_{t=1}^{T}w_{t}\eta_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$ $\leq\sum_{t=1}^{T}\sum_{s=1}^{t}\left(w_{s}-w_{s-1}\right)\eta_{t}\left(F(\boldsymbol{x}_{s})-F(\boldsymbol{x})\right)$
$=\sum_{t=1}^{T}\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\eta_{s}\right)\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right).$
Thus, we can lower bound the L.H.S. of (31) by
$\sum_{t=1}^{T}w_{t}\eta_{t}\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{y}_{t})\right)$ $=\sum_{t=1}^{T}w_{t}\eta_{t}\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right)-w_{t}\eta_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$
$\geq\sum_{t=1}^{T}\left[w_{t}\eta_{t}-\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\eta_{s}\right)\right]\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right)$
$=w_{T}\eta_{T}\left(F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right),$ (32)
where the last step is due to, for $t\in\left[T-1\right]$,
@@
$=\frac{\eta_{T}}{\sum_{s=t+1}^{T}\eta_{s}}\cdot\eta_{t}-\frac{\eta_{T}}{\sum_{s=t+1}^{T}\eta_{s}}\cdot\eta_{t}=0,$
@@
We plug (32) back into (31) and divide both sides by $w_{T}\eta_{T}$to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]$ $\leq\frac{w_{0}D^{2}}{2w_{T}\eta_{T}}+\sum_{t=1}^{T}\frac{w_{t}\eta_{t}^{2}}{w_{T}\eta_{T}}G^{2}+\mathsf{C}(\mathsf{p})\frac{w_{t}\eta_{t}^{\mathsf{p}}}{w_{T}\eta_{T}}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$

### v2

@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
[…]
As such, we can take expectations on both sides of (30) and rearrange terms to have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right]\leq\frac{(1-\mu\eta_{t})\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq$ $\frac{(1-\mu\eta_{t})\mathbb{\mathbb{E}}\left[\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}},$ (31)
where the second step is due to $\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\leq\left(1-\frac{w_{t-1}}{w_{t}}\right)\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t}\right\|^{2}+\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}=\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}$ by (27) and the convexity of $\left\|\boldsymbol{x}_{t}-\cdot\right\|^{2}$. Multiply both sides of (31) by $w_{t}\gamma_{t}$ and sum up from $t=1$ to $T$ to obtain (note that $\frac{\gamma_{t}}{\eta_{t}}=\frac{1}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\frac{\gamma_{t+1}(1-\mu\eta_{t+1})}{\eta_{t+1}},\forall t\in\left[T-1\right]$)
$\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right)\right]$
$\leq$ $\frac{\gamma_{1}(1-\mu\eta_{1})w_{0}\left\|\boldsymbol{x}_{1}-\boldsymbol{y}_{0}\right\|^{2}}{2\eta_{1}}-\frac{\gamma_{T}\mathbb{\mathbb{E}}\left[w_{T}\left\|\boldsymbol{x}_{T+1}-\boldsymbol{y}_{T}\right\|^{2}\right]}{2\eta_{T}}+\sum_{t=1}^{T}8w_{t}\gamma_{t}\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq$ $\frac{(1-\mu\eta_{1})w_{0}D^{2}}{2}+\sum_{t=1}^{T}8w_{t}\gamma_{t}\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}.$ (32)
@@
which implies that
$\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$ $\leq\sum_{t=1}^{T}\sum_{s=1}^{t}\left(w_{s}-w_{s-1}\right)\gamma_{t}\left(F(\boldsymbol{x}_{s})-F(\boldsymbol{x})\right)$
$=\sum_{t=1}^{T}\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\gamma_{s}\right)\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right).$
Thus, we can lower bound the L.H.S. of (32) by
$\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right)$ $=\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{x})\right)-w_{t}\gamma_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$
$\geq w_{T}\gamma_{T}\left(F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right)-\left(w_{1}-w_{0}\right)\left(\sum_{s=1}^{T}\gamma_{1}\right)\left(F(\boldsymbol{x}_{1})-F(\boldsymbol{x})\right)$
$+\sum_{t=2}^{T}\left[w_{t-1}\gamma_{t-1}-\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\gamma_{s}\right)\right]\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right)$
$=w_{T}\gamma_{T}\left(F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right),$ (33)
where the last step is due to, for $t\geq 2$,
[…]
$=\frac{\gamma_{T}}{\sum_{s=t-1}^{T}\gamma_{s}}\cdot\gamma_{t-1}-\frac{\gamma_{T}}{\sum_{s=t-1}^{T}\gamma_{s}}\cdot\gamma_{t-1}=0,$
[…]
We plug (33) back into (32) and divide both sides by $w_{T}\gamma_{T}$ to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right]$ $\leq\frac{(1-\mu\eta_{1})w_{0}D^{2}}{2w_{T}\gamma_{T}}+\sum_{t=1}^{T}\frac{8w_{t}\gamma_{t}\eta_{t}}{w_{T}\gamma_{T}}G^{2}+\frac{2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{w_{T}\gamma_{T}}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
[…]
Finally, relabel $T$ by T1T-1 to complete the proof. ∎
Equipped with Theorem 10, we show the following last-iterate convergence rate for $\mathsf{SGD}$$\mathsf{OGD}$. As far as we know, this is the first and the only provable result demonstrating that the last iterate of $\mathsf{SGD}$ can converge in heavy-tailed stochastic optimization without gradient clipping. When $T$ is unknown, we only incur an extra logarithmic factor compared to the best possible rate. In particular, the first stepsize degenerates to $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ employed in Theorem 1 if $H=0$. When $T$ is known, we employ the linear decay rate proposed by [50] to give an optimal rate.
Corollary 5.
Under Assumption 1 (without the third point) and additionally assuming (26) (with $\mu=0$) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$:
  • taking $\eta_{t}=\frac{1}{2H}\land\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta_{t}=\frac{T-t}{2HT}\land\frac{D(T-t)}{GT^{3/2}}\land\frac{D(T-t)}{\sigma T^{1+1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Since both kinds of stepsize satisfy $\eta_{t}\leq\frac{1}{2H\lor\mu}\overset{\mu=0}{=}\frac{1}{2H}$, by Theorem 10, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]$ $\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}}$
$=\frac{D^{2}}{\sum_{t=1}^{T-1}\eta_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T-1}\eta_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}}{\sum_{s=t}^{T-1}\eta_{s}},$ (34)
where the second step is due $\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\eta_{t}$ when $\mu=0$.
For $\eta_{t}=\frac{1}{2H}\land\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$, we observe that by Cauchy-Schwarz inequality
$(T-t)^{2}\leq\left(\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}\right)\left(\sum_{s=t}^{T-1}\eta_{s}\right)\Rightarrow\frac{1}{\sum_{s=t}^{T-1}\eta_{s}}\leq\frac{\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}.$
Thus, there is
[…]
We first bound
$\sum_{t=1}^{T-1}\frac{1}{\eta_{t}}=\sum_{t=1}^{T-1}2H\lor\frac{G\sqrt{t}}{D}\lor\frac{\sigma t^{1/\mathsf{p}}}{D}\leq\sum_{t=1}^{T-1}2H+\frac{G\sqrt{t}}{D}+\frac{\sigma t^{1/\mathsf{p}}}{D}\lesssim HT+\frac{G}{D}T^{3/2}+\frac{\sigma}{D}T^{1+1/\mathsf{p}},$
@@
$\frac{D^{2}}{(T-1)^{2}}\sum_{t=1}^{T-1}\frac{1}{\eta_{t}}\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$ (36)
Next, we know
$\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}$ $\overset{(a)}{\leq}\sum_{t=1}^{T-1}\left[\frac{2HD^{2}}{G^{2}}\cdot\frac{1}{t(T-t)}+\frac{D}{G}\cdot\frac{\sum_{s=t}^{T-1}\sqrt{s}}{t(T-t)^{2}}+\frac{\sigma D}{G^{2}}\cdot\frac{\sum_{s=t}^{T-1}s^{1/\mathsf{p}}}{t(T-t)^{2}}\right]$
@@
where $(a)$ is by $\eta_{t}\leq\frac{D}{G\sqrt{t}}$ and $\frac{1}{\eta_{s}}\leq 2H+\frac{G\sqrt{s}}{D}+\frac{\sigma s^{1/\mathsf{p}}}{D}$. Hence, there is
$G^{2}\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (37)
Similarly, we can bound
$\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (38)
Finally, we plug (36), (37) and (38) back into (35) to conclude.
For $\eta_{t}=\frac{T-t}{2HT}\land\frac{D(T-t)}{GT^{3/2}}\land\frac{D(T-t)}{\sigma T^{1+1/\mathsf{p}}}$, we denote by $\eta_{t}=\eta(T-t)$ for $\eta\triangleq\frac{1}{2HT}\land\frac{D}{GT^{3/2}}\land\frac{D}{\sigma T^{1+1/\mathsf{p}}}$ to have
@@
$=\frac{2D^{2}}{\eta T(T-1)}+2\eta G^{2}\sum_{t=1}^{T-1}\frac{T-t}{T-t+1}+2\eta^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{(T-t)^{\mathsf{p}-1}}{T-t+1}$
$\lesssim\frac{D^{2}}{\eta T^{2}}+\eta G^{2}T+\eta^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}T^{\mathsf{p}-1}\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
Next, Corollary 6 provides the first last-iterate convergence results in the strongly convex case. In particular, when $H=0\Rightarrow\kappa=0$, the first stepsize degenerates to $\eta_{t}=\frac{1}{\mu t}$, as used in Theorem 8, and yields a last-iterate rate $\frac{G^{2}\left(1+\log T\right)}{\mu T}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}T^{\mathsf{p}-1}}$, matching the average-iterate rate (up to an extra logarithmic factor) implied by the online-to-batch conversion for Theorem 8.
Corollary 6.
Under Assumption 1 (without the third point) and additionally assuming (26) (with $\mu>0$) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, let $\kappa\triangleq\frac{H}{\mu}$:
  • taking $\eta_{t}=\frac{1}{\mu(t+2\kappa)}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta_{t}=\begin{cases}\frac{1}{\mu(1+2\kappa)}&t\leq\tau\\ \frac{2}{\mu(t-\tau+2+4\kappa)}&t\geq\tau+1\end{cases}$ where $\tau\triangleq\left\lceil\frac{T}{2}\right\rceil$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{\exp\left(\frac{T}{2+4\kappa}\right)}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Since both kinds of stepsize satisfy $\eta_{t}\leq\frac{1}{2H\lor\mu}$, by Theorem 10, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}}.$ (39)
For $\eta_{t}=\frac{1}{\mu(t+2\kappa)}$, we can find
$\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\frac{1}{\mu(t+2\kappa)\prod_{s=2}^{t}\frac{s-1+2\kappa}{s+2\kappa}}=\frac{1}{\mu(1+2\kappa)}=\eta_{1}.$ (40)
Hence, there is